OpenStudy (anonymous):

find a third-degree polynomial equation with rational coefficients that has roots -4 and 2+i

OpenStudy (anonymous):

one factor is \((x+4)\) the harder part is finding the quadratic with zeros \(2+i\) and\(2-i\) but there are a couple quick ways to do it do you know any?

OpenStudy (anonymous):

no, i didn't understand this lesson when it was taught

OpenStudy (anonymous):

ok we can go slow

OpenStudy (anonymous):

since \(-4\) is a zero of the polynomial, one factor of the polynomial is \(x+4\) is that much ok?

OpenStudy (anonymous):

yea i get that

OpenStudy (anonymous):

ok good now you have a complex zero \(2+i\) is it clear that the other zero must be its conjugate \(2-i\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so what we need to do is find the quadratic (degree 2) polynomial with zeros of \(2+i\) and \(2-i\)

OpenStudy (anonymous):

the really annoying way, that you don't want to use, is to write this in factored form as \[(x-(2+i))(x-(2-i))\] which is not as hard as it seems but is a headache just the same there is a much easier way to recap, our goal is to find a quadratic \(x^2+bx+c\) whose zeros are \(2+i\) and \(2-i\)

OpenStudy (anonymous):

best way is to work backwards set \(x=2+i\) and see what you get steps are \[x=2+i\] subtract 2\[x-2=i\] square both sides (carefully) \[(x-2)^2=i^2=-1\] or \[x^2-4x+4=-1\] then add \(1\) to both sides and get \[x^2-4x+5\]

OpenStudy (anonymous):

this tells you \(x^2-4x+5\) is the polynomial with zeros \(2+i\) and \(2-i\) you can check if you like by using the quadratic formula and see that that is what you get there is also a really really snappy way that i can show you if you like

OpenStudy (anonymous):

yes please, if you don't mind

OpenStudy (anonymous):

it is snappy, but it requires memorizing something the quadratic with zeros \(a+bi\) and \(a-bi\) is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

in your case you have \(2+i\) so \(a=2,b=1\) please notice that there is no "\(i\)" in this calculation so \[x^2-2\times 2x+(2^2+1^2)\] or \[x^2-4x+5\] is your quadratic

OpenStudy (anonymous):

that is so much easier thank you!

OpenStudy (anonymous):

of course your final job is to multiply out \[(x+4)(x^2-4x+5)\] without making an algebra error i would cheat

OpenStudy (anonymous):

yw, but don't forget that method requires memorizing it, so if you don't you can't use it if you practice on 2 or 3 you will probably remember it quadratic with zeros \(3+2i\) is \[x^2-2\times 3x+(3^2+2^2)=x^2-6x+13\]

OpenStudy (anonymous):

ok